# Uncertainty, von Neumann entropy and squeezing in bipartite state of two-level atoms

### v1

@@ Abstract
pacs:
03.67.Mn, 03.65.Ud, 42.50.Dv, 42.50.-p
@@ I I INTRODUCTION
Uncertainties in the simultaneous measurements of incompatible observables of quantum systems make quantum systems drastically different from classical ones. Uncertainty principle, first formulated by Heisenberg Heisenberg forms a central part of our understanding of quantum mechanics. Many works have been done on further developments and applications of uncertainty principle with the advent and progress of quantum information theory Robertson -Horodecki1 .
@@
Spin squeezing is an important experimental tool that have applications in the detection of quantum entanglement, high-precision metrology, atomic clocks, quantum Fisher information, quantum chaos, quantum phase transitions, Bose-Einstein condensates, and ion-trap experiments in quantum optics Jian .
@@
In this paper we consider a general entangled state of two two-level atoms. We show that though the von Neumann entropies of the entangled subsystems are always greater than that of the system as a whole, thus violating the entropic inequalities Horodecki1 , but the uncertainties of the subsystems are less than the system as a whole for certain range of the superposing constants of the quantum state. This is in contradiction with the consensus that more is the entropy, more is the uncertainty and more is the disorder. We also establish analytically a general relationship between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the system as a whole. Now, since the spin squeezing and spectroscopic squeezing are experimentally measurable, so, this provides an operational measure of the entanglement entropy of the subsystems of two-level atoms.
@@
In section II we briefly discuss the two-level atoms and pseudo-spin operators. In section III we calculate the uncertainties of the pseudo-spin operators in two mutually orthogonal directions in a plane perpendicular to the mean spin vector of the general entangled state of two two-level atoms, both for the composite state and the subsystems. We show that though the von Neumann entropy of the composite state is less than those of the subsystems, but, the uncertainties of the composite state are greater than those of the subsystems for certain range of the superposing constants of the quantum state. In section IV we establish general relationships between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing and spectroscopic squeezing parameters of the composite state. In section V we present the conclusion.
I.1 II TWO-LEVEL ATOM, PSEUDO-SPIN OPERATORS
@@
An atom has many electronic energy levels, but when it interacts with an external monochromatic electromagnetic radiation, the atom makes a transition from one of its energy level to the other. In this case, we mainly concentrate on those two energy levels and hence the atom is called as a two-level atom. Now, if among the assembly of $N$ such two-level atoms, the $n$-th atom has the upper and lower energy levels, denoted as $|u_{n}\rangle$ and $|l_{n}\rangle$, respectively, then, we can construct the pseudo-spin operators (with $\hbar=1$),
$\hat{J}_{n_{x}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle l_{n}|+|l_{n}\rangle\langle u_{n}|\big{)},$ (1)
$\hat{J}_{n_{y}}$ $=$ $(-i/2)\big{(}|u_{n}\rangle\langle l_{n}|-|l_{n}\rangle\langle u_{n}|\big{)},$ (2)
$\hat{J}_{n_{z}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle u_{n}|-|l_{n}\rangle\langle l_{n}|\big{)},$ (3)
@@
$[\hat{J}_{n_{x}},\hat{J}_{n_{y}}]=i\hat{J}_{n_{z}},$ (4)
and two more relations with cyclic changes in $x$, $y$ and $z$ Itano . For the entire system of two two-level atoms, we have collective pseudo-spin operators,
$\hat{J}_{x}=\sum_{i=1}^{2}\hat{J}_{i_{x}},\hat{J}_{y}=\sum_{i=1}^{2}\hat{J}_{i_{y}},\hat{J}_{z}=\sum_{i=1}^{2}\hat{J}_{i_{z}},$ (5)
where each term in the above summations is multiplied by the identity operators for all of the other atoms Itano , Sakurai .
@@
$\big{[}\hat{J}_{2_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $i\hat{J}_{2_{z}},...$ (6)
where the subscripts $1$ and $2$ are for atoms $1$ and $2$ respectively.
As a consequence of these commutation relations, the collective pseudo-spin operators $\hat{J}_{x}$, $\hat{J}_{y}$ and $\hat{J}_{z}$ satisfy,
$[\hat{J}_{x},\hat{J}_{y}]=i\hat{J}_{z}$ (7)
and two more relations with cyclic changes in $x$, $y$ and $z$.
The eigenvectors of $\hat{J}_{n_{z}}$ are denoted as $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively.
The use of these collective spin operators as discussed in Eqs. (5) are widely used in the problems of quantum optics Korbicz1 -Miyake2 .
@@ I.2 III UNCERTAINTIES AND VON NEUMANN ENTROPY IN A SYSTEM OF TWO TWO-LEVEL ATOMS
We, now, consider a system of two entangled two-level atoms, say atoms $a$ and $b$. A general pure state of this system in Schmidt decomposition is given as
$|\psi_{ab}\rangle=\sum_{i=1}^{2}c_{i}|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle,$ (8)
where $|\psi_{ab}\rangle$in $H_{a}\otimes H_{b}$, $c_{i}$ are the Schmidt coefficients and $\{|e^{i}_{a}\rangle\otimes|e^{i}_{b}\rangle\}$ is a orthonormal product basis. The Schmidt coefficients $c_{i}$ satisfy
@@
For the system of two two-level atoms we take the orthonormal product basis, made up of generalized superpositions of $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, as
@@
where $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ are the eigenvectors of $\hat{J}_{a_{z}}$ and $\hat{J}_{b_{z}}$ with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$ respectively and for the sake of simplicity $c_{3}$, $c_{4}$, $c_{5}$ and $c_{6}$ are assumed to be real. Since the basis vectors are normalized, we have
${c_{3}}^{2}+{c_{4}}^{2}$ $=$ $1,$ (10)
${c_{5}}^{2}+{c_{6}}^{2}$ $=$ $1$ (11)
@@
$\otimes\big{(}c_{6}\Big{|}\frac{1}{2}\Big{\rangle}-c_{5}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}$ (12)
@@
$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}$ (13)
For making the calculations simple we let
@@
In order to study the spin squeezing and spectroscopic squeezing properties of the system we need to calculate the coordinate frame independent inherent uncertainties in the components of $\hat{\mathbf{J}}$ in two mutually orthogonal directions, in a plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ of the system Kitagawa ; Wineland2 , where
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{n_{1}}\rangle\hat{n}_{1}+\langle\hat{J}_{n_{2}}\rangle\hat{n}_{2}+\langle\hat{J}_{n_{3}}\rangle\hat{n}_{3}.$ (19)
$\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ are three mutually orthogonal unit vectors. The expectation values as mentioned above are to be calculated over the quantum state in Eq. (12).
Now, to understand it concretely we take the directions of $\hat{n}_{1}$, $\hat{n}_{2}$ and $\hat{n}_{3}$ along the positive $x$, $y$ and $z$ axes respectively so that $\langle\hat{\mathbf{J}}\rangle$ is
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{x}\rangle\hat{i}+\langle\hat{J}_{y}\rangle\hat{j}+\langle\hat{J}_{z}\rangle\hat{k},$ (20)
and hence the uncertainties are the standard deviations
$\Delta J_{x,y}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x,y}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x,y}|\psi_{ab}\rangle^{2}}.$ (21)
Now, since for an arbitrary quantum state, as that given in Eq. (12) the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ points in an arbitrary direction in space, we conventionally rotate the coordinate system $\{x,y,z\}$ to $\{x^{\prime},y^{\prime},z^{\prime}\}$, such that $\langle\hat{\bf{J}}\rangle$ points along the $z^{\prime}$ axis and calculate the standard deviations of $\hat{J}_{x^{\prime}}$ and $\hat{J}_{y^{\prime}}$ for the state $|\psi_{ab}\rangle$. These are
$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle-\langle\psi_{ab}|\hat{J}_{x^{\prime},y^{\prime}}|\psi_{ab}\rangle^{2}}.$ (22)
Now, since the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ after rotation of the coordinate frame points along the $z^{\prime}$ axis, we have
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle=\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle=0,$ (23)
and hence, the standard deviations reduce to
$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi_{ab}|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi_{ab}\rangle}.$ (24)
We now proceed to calculate $\Delta J_{x^{\prime},y^{\prime}}$. We find that
$\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$ $=$ $(A+D)(B+C)$ (25)
@@
$\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$ $=$ $A^{2}-D^{2}.$ (27)
Therefore, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ lies in the $x-z$ plane. To align it along the $z^{\prime}$ direction we perform the following rotation.
@@
$\hat{J}_{z^{\prime}}$ $=$ $\hat{J}_{x}\sin\theta+\hat{J}_{z}\cos\theta,$ (30)
where,
@@
It can be verified using Eqs. (26), (28), (29) and (31) that
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle$ $=$ 00 0 (32)
$\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$ (33)
and hence the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ is now along the $z^{\prime}$-axis.
Now, using Eqs. (28) and (31), we get

### v2

# Uncertainty, von Neumann Entropy, and Squeezing in a Bipartite State of Two-Level Atoms

@@ Abstract
keywords: uncertainty, von Neumann entropy, two-level atom, spin squeezing, spectroscopic squeezing
@@ I I INTRODUCTION
Uncertainties in the simultaneous measurements of incompatible observables of quantum systems, make quantum systems drastically different from classical ones. Uncertainty principle, first formulated by Heisenberg Heisenberg , forms a central part of our understanding of quantum mechanics. Many works have been done on further developments and applications of uncertainty principle, with the advent and progress of quantum information theory Robertson -Horodecki1 .
@@
Spin squeezing is an important experimental tool that has potential applications in the detection of quantum entanglement, high-precision metrology, atomic clocks, quantum Fisher information, quantum chaos, quantum phase transitions, Bose-Einstein condensates, and ion-trap experiments in quantum optics. Spin squeezing is the reduction of quantum fluctuations in collective spin observables of atomic systems, typically measured by the variance of the spin components. This is achieved by correlations between the spins, which can be generated through nonlinear interactions or entanglement Kitagawa ; Jian .
Spectroscopic squeezing, on the other hand, refers to the reduction of quantum fluctuations in the spectral response of an atomic ensemble. This is typically measured by the variance of the absorption or emission spectrum. It has significant applications in enhancing resolutions and precisions in optical measurements, which lead to more accurate determination of atomic and molecular energy levels. Squeezing of spectroscopic signals enables the detection of weaker signals and more precise characterization of atomic and molecular systems. It enables the achievement of quantum-limited spectroscopy, where the measurement precision is limited only by the fundamental laws of quantum mechanics. It helps to develop more accurate atomic clocks, which rely on precise spectroscopic measurements. In quantum information processing, spectroscopic squeezing enables the development of more precise quantum gates and quantum computing architectures Wineland2 ; Jian .
@@
In this paper, we consider a general entangled state of two two-level atoms. We show that, although the von Neumann entropies of the entangled subsystems are always greater than that of the system as a whole —thus violating the entropic inequalities Horodecki1 —the uncertainties of the subsystems are lower than those of the system as a whole for certain ranges of the superposition constants of the quantum state. This is in contradiction with the idea that, the more is the entropy, more is the uncertainty, and the more is the disorder. We also establish, analytically, generalized relationships between the von Neumann entropies of the subsystems, and the uncertainties, spin squeezing, and spectroscopic squeezing parameters of the system as a whole. Now, since spin squeezing and spectroscopic squeezing are experimentally measurable, this provides an operational measure of the entanglement entropy of the subsystems of two-level atoms.
It is worthwhile to mention here that noise and decoherence are significant practical challenges in experiments involving spin and spectroscopic squeezing. Noise and decoherence can destroy spin coherence, introduce spin fluctuations, and limit the squeezing factors in spin squeezing experiments. In spectroscopic squeezing experiments, they can broaden spectral lines, introduce spectral fluctuations, and limit squeezing factors. However, a lot of experimental and theoretical progress has been made in the study of spin and spectroscopic squeezing in many physical systems —for example, in ion traps and Bose —Einstein condensates [33 - 58].
@@
In section II, we briefly discuss two-level atoms, pseudo-spin operators, and spin and spectroscopic squeezing. In section III, we calculate the uncertainties of the pseudo-spin operators in two mutually orthogonal directions, in a plane perpendicular to the mean spin vector of the general entangled state of two two-level atoms, both for the composite state and the subsystems. We show that, although the von Neumann entropy of the composite state is less than those of the subsystems, the uncertainties of the composite state are greater than those of the subsystems for certain ranges of the superposing constants of the quantum state. In section IV, we establish general relationships between the von Neumann entropies of the subsystems and the uncertainties, spin squeezing, and spectroscopic squeezing parameters of the composite state. In section V, we present the conclusion.
I.1 II TWO-LEVEL ATOM, PSEUDO-SPIN OPERATORS, SPIN AND SPECTROSCOPIC SQUEEZING
@@
An atom has many electronic energy levels, but when it interacts with external monochromatic electromagnetic radiation, it undergoes a transition from one energy level to another. In this case, we mainly focus on those two energy levels; hence, the atom is referred to as a two-level atom. Now, if among an assembly of $N$ such two-level atoms, the $n$-th atom has upper and lower energy levels denoted by $|u_{n}\rangle$ and $|l_{n}\rangle$, respectively, then we can construct the pseudo-spin operators (with $\hbar=1$),
$\hat{J}_{n_{x}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle l_{n}|+|l_{n}\rangle\langle u_{n}|\big{)},$
$\hat{J}_{n_{y}}$ $=$ $(-i/2)\big{(}|u_{n}\rangle\langle l_{n}|-|l_{n}\rangle\langle u_{n}|\big{)},$
$\hat{J}_{n_{z}}$ $=$ $(1/2)\big{(}|u_{n}\rangle\langle u_{n}|-|l_{n}\rangle\langle l_{n}|\big{)},$
@@
$[\hat{J}_{n_{x}},\hat{J}_{n_{y}}]=i\hat{J}_{n_{z}},$
and two more relations with cyclic changes in $x$, $y$ and $z$ Itano . Since these operators are not actual spin angular momentum operators, but satisfy the same commutation relations as those of spin-$\frac{1}{2}$ particles, they are referred to as pseudo-spin operators. For the entire system of two two-level atoms, we define the collective pseudo-spin operators
$\hat{J}_{x}=\sum_{i=1}^{2}\hat{J}_{i_{x}},\hat{J}_{y}=\sum_{i=1}^{2}\hat{J}_{i_{y}},~{}~{}and~{}~{}\hat{J}_{z}=\sum_{i=1}^{2}\hat{J}_{i_{z}},$
where each term in the summations is implicitly tensor-multiplied by identity operators acting on the other atom Itano , Sakurai .
@@
$\big{[}\hat{J}_{2_{x}},\hat{J}_{2_{y}}\big{]}$ $=$ $i\hat{J}_{2_{z}},...,$
where the subscripts $1$ and $2$ are for atoms $1$ and $2$, respectively.
As a consequence of these commutation relations, the collective pseudo-spin operators $\hat{J}_{x}$, $\hat{J}_{y}$, and $\hat{J}_{z}$ satisfy
$[\hat{J}_{x},\hat{J}_{y}]=i\hat{J}_{z},$ (1)
and two more relations with cyclic changes in $x$, $y$, and $z$.
The eigenvectors of $\hat{J}_{n_{z}}$ are denoted as $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, with eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$, respectively.
The use of these collective spin operators, as discussed in Eq. (I.1), is widely used in problems of quantum optics Korbicz1 -Miyake2 .
In order to study the spin squeezing and spectroscopic squeezing properties of a quantum state, say $|\psi\rangle$, of a system of two-level atoms, we need to calculate the coordinate-frame independent inherent uncertainties in the components of $\hat{\mathbf{J}}$ along two mutually orthogonal directions in a plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ of the system Kitagawa ; Wineland2 , where
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{n_{1}}\rangle\hat{n}_{1}+\langle\hat{J}_{n_{2}}\rangle\hat{n}_{2}+\langle\hat{J}_{n_{3}}\rangle\hat{n}_{3}.$
$\hat{n}_{1}$, $\hat{n}_{2}$, and $\hat{n}_{3}$ are three mutually orthogonal unit vectors. The expectation values mentioned above are to be calculated with respect to the quantum state $|\psi\rangle$.
To understand this concretely, we take the directions of $\hat{n}_{1}$, $\hat{n}_{2}$, and $\hat{n}_{3}$ along the positive $x$, $y$, and $z$ axes, respectively, so that
$\langle\hat{\mathbf{J}}\rangle=\langle\hat{J}_{x}\rangle\hat{i}+\langle\hat{J}_{y}\rangle\hat{j}+\langle\hat{J}_{z}\rangle\hat{k},$
and hence the uncertainties are given by the standard deviations
$\Delta J_{x,y}=\sqrt{\langle\psi|{\hat{J}_{x,y}}^{2}|\psi\rangle-\langle\psi|\hat{J}_{x,y}|\psi\rangle^{2}}.$ (2)
Now, since, for an arbitrary quantum state $|\psi\rangle$, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ points in an arbitrary direction in space, we conventionally rotate the coordinate system $\{x,y,z\}$ to $\{x^{\prime},y^{\prime},z^{\prime}\}$, such that $\langle\hat{\bf{J}}\rangle$ points along the $z^{\prime}$ axis, and calculate the standard deviations of $\hat{J}_{x^{\prime}}$ and $\hat{J}_{y^{\prime}}$ for the state $|\psi\rangle$. Since the mean spin vector $\langle\hat{\mathbf{J}}\rangle$, after rotation of the coordinate frame, points along the $z^{\prime}$ axis, we have
$\langle\psi|\hat{J}_{x^{\prime}}|\psi\rangle=\langle\psi|\hat{J}_{y^{\prime}}|\psi\rangle=0,$
and hence the standard deviations, as mentioned in Eq. (2), reduce to
$\Delta J_{x^{\prime},y^{\prime}}=\sqrt{\langle\psi|{\hat{J}_{x^{\prime},y^{\prime}}}^{2}|\psi\rangle}.$ (3)
In this paper, we calculate these coordinate-independent uncertainties for a general pure state of two two-level atoms, both for the composite state and for the individual entangled parts.
For a system of $N$ two-level atoms, the spin squeezing parameter, introduced by Kitagawa et al. Kitagawa ; Jian , is defined as
$\xi_{s}=\sqrt{\frac{2}{j}}\Delta J_{\overrightarrow{n}_{\perp}},$
where $\overrightarrow{n}_{\perp}$ lies in the plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$, and $j=\frac{N}{2}$. Now, conventionally, we take $\overrightarrow{n}_{\perp}$ as the directions of $x^{\prime}$ and $y^{\prime}$ axes, and hence, the spin squeezing parameters become
$\xi_{s_{x}}=\sqrt{\frac{2}{j}}\Delta J_{x^{\prime}}~{}~{}and~{}~{}\xi_{s_{y}}=\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}.$ (4)
If, for a quantum state, $\xi_{s_{x}}<1$, then we have $\xi_{s_{y}}>1$, respecting Heisenberg’s uncertainty principle, and we conclude that we have spin squeezing in the $x$-quadrature —and vice versa. The degree of spin squeezing is quantified by this spin squeezing parameter. The smaller $\xi_{s}$ is (i.e., the more $\xi_{s}<1$), the greater the spin squeezing.
The spectroscopic squeezing parameter $\xi_{R}$, introduced by Wineland et al. Wineland2 ; Jian , is given by
$\xi_{R}=\sqrt{N}\frac{\Delta J_{\overrightarrow{n}_{\perp}}}{|\langle\hat{\mathbf{J}}\rangle|}.$
In a similar fashion, as mentioned above, we take the directions of $\overrightarrow{n}_{\perp}$ as the directions of $x^{\prime}$ and $y^{\prime}$ axes, and hence, the spectroscopic squeezing parameters become
$\xi_{R_{x}}=\sqrt{N}\frac{\Delta J_{x^{\prime}}}{|\langle\hat{\mathbf{J}}\rangle|}~{}~{}and~{}~{}\xi_{R_{y}}=\sqrt{N}\frac{\Delta J_{y^{\prime}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (5)
If $\xi_{R_{x}}$ or $\xi_{R_{y}}$ $<1$, we have spectroscopic squeezing in the system. Spectroscopic squeezing is quantified using these parameters. The smaller the value of $\xi_{R}<1$, the greater the spectroscopic squeezing.
We now proceed to study the coordinate-independent inherent uncertainties and the von Neumann entropy of a general system of two two-level atoms.
@@ I.2 III UNCERTAINTIES AND VON NEUMANN ENTROPY IN A SYSTEM OF TWO TWO-LEVEL ATOMS
A general pure state of two systems, say $A$ and $B$, is given by
$|\psi_{AB}\rangle=\sum_{i=1}^{2}c_{i}|e^{i}_{A}\rangle\otimes|e^{i}_{B}\rangle,$
where $|\psi_{AB}\rangle$in $H_{A}\otimes H_{B}$, $c_{i}$ are the Schmidt coefficients, and $\{|e^{i}_{A}\rangle\otimes|e^{i}_{B}\rangle\}$ form an orthonormal product basis. The Schmidt coefficients $c_{i}$ satisfy
@@
We now consider a system of two entangled two-level atoms, denoted as atoms $a$ and $b$. To construct the composite quantum pure state $|\psi_{ab}\rangle$ of the two atoms, we take the orthonormal product basis, composed of generalized superpositions of $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$, given by
@@
where $|\frac{1}{2}\rangle$ and $|-\frac{1}{2}\rangle$ are the eigenvectors of $\hat{J}_{a_{z}}$ and $\hat{J}_{b_{z}}$, corresponding to eigenvalues $\frac{1}{2}$ and $-\frac{1}{2}$, respectively. For the sake of simplicity, $c_{3}$, $c_{4}$, $c_{5}$, and $c_{6}$ are assumed to be real. Since the basis vectors are normalized, we have
${c_{3}}^{2}+{c_{4}}^{2}=1$
and
${c_{5}}^{2}+{c_{6}}^{2}=1.$
@@
$\otimes\big{(}c_{6}\Big{|}\frac{1}{2}\Big{\rangle}-c_{5}\Big{|}-\frac{1}{2}\Big{\rangle}\big{)}\bigg{]}.$
@@
$\Big{(}c_{1}c_{4}c_{6}+c_{2}c_{3}c_{5}\Big{)}\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}.$ (7)
To simplify the calculations, we let
@@
Therefore, Eq. (7) becomes
$|\psi_{ab}\rangle=A\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+B\Big{|}\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}+$
$C\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}\frac{1}{2}\Big{\rangle}+D\Big{|}-\frac{1}{2}\Big{\rangle}\otimes\Big{|}-\frac{1}{2}\Big{\rangle}.$ (13)
We now proceed to calculate the coordinate independent inherent uncertainties, $\Delta J_{x^{\prime},y^{\prime}}$, as mentioned in the previous section. To that end, we first calculate $\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$, $\langle\psi_{ab}|\hat{J}_{y}|\psi_{ab}\rangle$, and $\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$. We find that
$\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle$ $=$ $(A+D)(B+C),$ (14)
@@
$\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle$ $=$ $A^{2}-D^{2},$ (16)
and hence, the square of the magnitude of the mean spin vector is
$|\langle\hat{\mathbf{J}}\rangle|^{2}=\langle\psi_{ab}|\hat{J}_{x}|\psi_{ab}\rangle^{2}+\langle\psi_{ab}|\hat{J}_{z}|\psi_{ab}\rangle^{2}$
$=A^{2}(A^{2}+B^{2}+C^{2}+2BC)+$
$D^{2}(B^{2}+C^{2}+D^{2}-2A^{2}+2BC)+$
$2AD(B+C)^{2}.$ (17)
We notice from Eqs. (14) to (16) that the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ lies in the $x-z$ plane. To align it along the $z^{\prime}$ direction, we perform the following rotation:
@@
$\hat{J}_{z^{\prime}}$ $=$ $\hat{J}_{x}\sin\theta+\hat{J}_{z}\cos\theta,$
where
@@
It can be verified using Eqs. (15), (18), (19), and (20) that
$\langle\psi_{ab}|\hat{J}_{x^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$
$\langle\psi_{ab}|\hat{J}_{y^{\prime}}|\psi_{ab}\rangle$ $=$ $0,$
and hence, the mean spin vector $\langle\hat{\mathbf{J}}\rangle$ is now aligned along the $z^{\prime}$-axis.
Using Eqs. (18) and (20), we obtain